OpenStudy (ammarah):

Ppl who have a geometry textbook. Go to page 91 #14

OpenStudy (anonymous):

there are 200 Geo textbooks across the world, likely more. we will need more info

OpenStudy (ammarah):

The California geometry book

OpenStudy (ammarah):

Pg 100 #18

OpenStudy (ammarah):

Wats 90 divided by 15?

OpenStudy (anonymous):

90/15= 18/3=6.

OpenStudy (anonymous):

there are 20 California ones,

OpenStudy (ammarah):

Oh